OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 4x^2+6x+2=0 Numerical Answers Expected!

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

b^2-4(a*c) Answers: -1, & -2

OpenStudy (anonymous):

explain please?

OpenStudy (anonymous):

actually i'm sorry the answers are -1/2 & -1 my bad and when you use the quadratic formula: x= -b+ or - square root b^2- 4(a*c) all over 2*a. you may want to check YouTube for the Quadratic equation song to help you better understand it but it's simply plugging in the numbers into the formula.

OpenStudy (anonymous):

when you take the square root of the number you're able to identify if the number is irrational or rational meaning if it has a decimal or not and in this case, it's a whole number which means it's rational

OpenStudy (anonymous):

That's not why. If the stuff under the square root (the discriminant) is negative, then you don't have any real solutions and can't factor as a result. You can certainly factor if your solutions are irrational, it just might not be pleasant or feasible to do on an exam. For example, we can factor \(x^2-2\) into \((x-\sqrt{2})(x+\sqrt{2})\) but we can't factor \(x^2+1\) because it has only complex solutions (namely \(\sqrt{-1}\) and \(-\sqrt{-1}\) ).

OpenStudy (anonymous):

im confused? what's the answer??????

OpenStudy (anonymous):

you can factor precisely when \(b^2-4ac\) is non-negative, but it may be difficult if \(b^2-4ac\) isn't rational. Soccergurl gave you the solutions for the second part.

OpenStudy (anonymous):

sooo the answers are -1/2 & -1

OpenStudy (anonymous):

Looks like it.

OpenStudy (anonymous):

are you sure my answer is supposed to have two numbers in it?

OpenStudy (anonymous):

If \(b^2-4ac\) is zero, then you have 1 solution. If \(b^2-4ac\) is positive, then you have 2 solutions. If \(b^2-4ac\) is negative, then you have 0 solutions.

OpenStudy (anonymous):

ok